OpenStudy (anonymous):

A WHEEL MAKES 10 3/4 REVOLUTIONS PER MINUTE. IF IT ROTATES FOR 34 MINUTES, HOW MANY REVOLUTIONS DOES IT MAKE?

OpenStudy (anonymous):

ten and three quarters times thirty four

OpenStudy (skullpatrol):

@RUBY91 does that make sense?

OpenStudy (skullpatrol):

$$\Huge 10\frac{3}{4}\frac{revolutions}{minute}*34minutes$$

OpenStudy (anonymous):

365.5 ??BUT I HAVE TO TURN THAT INTO A PROPER FRACTION OR MIXED NUMBER

OpenStudy (skullpatrol):

0.5 = 1/2